OpenStudy (anonymous):

find the distinguisable permutations in the group of letters: C A L C U L U S

OpenStudy (anonymous):

You're trying to find all possible combinations of the letters, without distinguishing between 'different' U's, C's and L's. If you have 8 letters (as here), the total number of permutations taken 8 at a time is 8!. But, you have C, L and U's counted 2! times each (you can arrange C, 2! ways, for example). You must undo this excess counting, so you have\[\frac{8!}{2!^3}=\frac{8 \times 7!}{8}=7!=5040\]distinguishable permutations.

OpenStudy (anonymous):

so after i work out the problem ill have my answr? or do i leave it as is

OpenStudy (anonymous):

Oh, I think you're being asked to find the number of distinguishable permutations. It would take you forever to write them all out (there's 5040 of them!).

OpenStudy (anonymous):

so the answers 5040? (:

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

im sorry! im terrible at math, im more of a science gal, but thank you soo much!

OpenStudy (anonymous):

no probs :)